find the nth term of the arithmetic sequence described,
-7,-3,1,5...for n = 35

Answer by mmm4444bot(95)

Hello There:
-7, -3, 1, 5, ...
We see that each element is four more than the one before, so the common difference between each pair of elements is four.
This is an arithmetic sequence.
The general form is:
nth term = first term + common difference * (n - 1)
The first term is -7.
The common difference is 4.
So, we have nth term = -7 + 4*(n - 1)
The 35th term is: -7 + 4(35 - 1) = -7 + 4(34) = -7 + 136 = 129
~ Mark